What will be the equivalent weight of Fe2S3 in the reactionFe2S3 ----> 2Fe(2+) + SO2

Grade:11

2 Answers

Vikas TU
7 years ago
Eq. Weight = Toatal Molar mass/Valence factor. => 208/v.f
V.f. for Fe here calculating as:
=> 2*(3-2) = > 2
Hence Eq. weight = 208/2 = > 104.
Sumeet Gupta
5 years ago
Hi , to solve this problem for these steps
 
A)Find n factor for the reaction.
n =2×3+6×2+1×4 =22
B)Now equivalent weight is mol weight /n factor
M/22 =208/22 =9.4545 gms